1.            A bullet moving at the rate of 300 m/s is fired into a thick target and penetrates up to 500 mm. If it is fired into a 250 mm thick target, find the velocity of emergence. Take the resistance to be uniform in both the cases.



1
Expert's answer
2021-07-01T09:34:20-0400

The given information is

  • The initial velocity is "V_{(o)}=300\;\text{m}/\text{s}"
  • The initial distance traveled is "d_{(i)}=500\;\text{mm}"
  • The final distance traveled is "d_{(f)}=250\;\text{mm}"

Converting the distances to meters.

Equivalence millimeters to meters. "1\;\text{m}=1000\;\text{mm}"

Multiplying by the conversion factor.

"d_{(i)}=500\;\text{mm}\times \dfrac{1\;\text{m}}{1000\;\text{mm}}=0.5\;\text{m}\\
d_{(f)}=250\;\text{mm}\times \dfrac{1\;\text{m}}{1000\;\text{mm}}=0.25\;\text{m}\\"


The acceleration that the bullet undergoes in the first impact is

"a=\dfrac{V_{(f)}^{2}-V_{(i)}^{2}}{2\;d_{(i)}}"

In which.

"V_{(i)}" is the initial velocity.

"V_{(f)}" is the final velocity.

"d_{(i)}" is the distance


Evaluating numerically.

"a=\dfrac{V_{(f)}^{2}-V_{(i)}^{2}}{2\;d_{(i)}}\\
a=\dfrac{(0\;\text{m}/\text{s})^{2}-(300\;\text{m}/\text{s}^{2}}{2\times 0.5\;\text{m}}\\
a=-90,000\;\text{m}/\text{s}^{2}"


Now, for the second shot, the final velocity is given by

"V_{(f)}^{2}=V_{(i)}^{2}+2\;a\;d_{(f)}"

\\

Where.

In which.

"V_{(i)}" is the initial velocity.

"V_{(f)}" is the final velocity.

"d_{(f)}" is the distance


Evaluating numerically.

"V_{(f)}^{2}=V_{(i)}^{2}+2\;a\;d_{(f)}\\
V_{(f)}=\sqrt{V_{(i)}^{2}+2\;a\;d_{(f)}}\\
V_{(f)}=\sqrt{(300\;\text{m}/\text{s})^{2}+2\times -90,000;\text{m}/\text{s}^{2}\times 0.25\;\text{m}}\\
V_{(f)}=\sqrt{ 45,000\text{m}^{2}/\text{s}^{2}}\\"

"V_{(f)}=212\;\text{m}/\text{s}"


Answer.

The velocity with which it emerges from the second target is "\displaystyle \color{red}{\boxed{V_{(f)}=212\;\text{m}/\text{s}}}"